# arXiv:2603.02378v1[cs.CR] 02 Mar 2026

### v1

@@ Abstract
Cryptographic provenance standards such as C2PA and invisible watermarking are positioned as complementary defenses for content authentication, yet the two verification layers are technically independent: neither conditions on the output of the other. This work formalizes and empirically demonstrates the Integrity Clash, a condition in which a digital asset carries a cryptographically valid C2PA manifest asserting human authorship while its pixels simultaneously carry a watermark identifying it as AI-generated, with both signals passing their respective verification checks in isolation. We construct metadata washing workflows that produce these authenticated fakes through standard editing pipelines, requiring no cryptographic compromise, only the semantic omission of a single assertion field permitted by the current C2PA specification. To close this gap, we propose a cross-layer audit protocol that jointly evaluates provenance metadata and watermark detection status, achieving 100% classification accuracy across 3,500 test images spanning four conflict-matrix states and three realistic perturbation conditions. Our results demonstrate that the gap between these verification layers is unnecessary and technically straightforward to close.
@@ 1 Introduction
This framing, however, obscures a structural vulnerability that follows directly from the architecture of both systems. The two verification layers are technically independent where C2PA manifest validation is a cryptographic signature check over metadata [8, 7, 45], while watermark detection is a signal-decoding operation over pixel data [54, 55]. Neither procedure conditions on the output of the other. As a result, a digital asset can carry a cryptographically valid C2PA manifest asserting one provenance history while its pixels simultaneously carry a watermark signal that contradicts that history, with both signals passing their respective verification checks in isolation. In this paper, we formalize and empirically demonstrate this condition, which we refer to as the Integrity Clash. Concretely, an AI-generated image can be passed through a C2PA-compliant editing tool, signed with valid credentials that assert only a human editing action, and emerge with a valid provenance manifest while still containing a watermark that identifies it as synthetically generated. Any verifier that inspects only one layer receives an incomplete account of the asset’s origin.
@@
Prior work has examined each verification layer in isolation. Research on watermarking evaluates embedding robustness under compression, cropping, and adversarial attack [4, 53, 43], while research on provenance standards analyzes manifest integrity, soft-binding collision risks, and metadata stripping vulnerabilities [7, 9]. Recent empirical audits have documented fragmented adoption of these systems across generative AI platforms [37, 42], and preliminary frameworks have proposed multi-signal consistency checks as a forensic methodology [45, 23]. However, no existing work empirically demonstrates the cross-layer contradiction as a concrete, reproducible attack workflow, nor provides a systematic protocol for detecting such contradictions in deployed pipelines. This paper addresses that gap. We make three contributions:
@@ 2.2 Invisible Watermarking for Generative AI
Recent work has shown that invisible watermarks are not immutable. Benchmarks such as WAVES [4] systematically evaluate robustness under compression, rescaling, and noise, and adversarial studies have demonstrated practical removal and transplantation attacks against open-source watermarking tools [31, 35]. The winning system in the NeurIPS 2024 Invisible Watermark Removal Challenge [44, 15], for example, uses generative-model-based regeneration to strip a variety of watermark signals while maintaining high perceptual quality, and Zhao et al. [53] prove that a broad class of invisible image watermarks are removable using generative AI. These findings establish that watermark signals can be weakened under determined attack, but also confirm that under standard editorial workflows, watermarks remain persistent to produce detectable cross-layer inconsistencies with co-existing metadata.
@@ 2.3 Related Work
Ecosystem audits have begun to document the gap between these layers. Rijsbosch et al. [42] report that only 38% of surveyed AI image generators implement adequate watermarking and just 18% integrate with C2PA-style labeling standards, establishing that the fragmentation enabling cross-layer inconsistencies is empirically widespread. Nemecek et al. [37] argue that without independent verification across technical, audit, and enforcement layers, watermarking serves as symbolic compliance rather than effective oversight. Simmons and Winograd [45] define success and failure scenarios when provenance and watermark signals co-exist in broadcast media, while agent-based forensic pipelines have been explored for multimodal detection [27]. These contributions identify the fragmentation problem or propose frameworks, but none empirically demonstrate the cross-layer contradiction as a reproducible attack workflow or evaluate a detection protocol against it.
Industry and standards-focused analyses reinforce this picture at a systems level. Recent work on media integrity [51] concludes that high-confidence validation requires aligning multiple signals rather than trusting any single mechanism in isolation, and the Origin Lens framework [30] advocates combined provenance verification and AI detection but does not examine cases where a valid manifest coexists with a contradictory watermark. None, to our knowledge, empirically demonstrates the cross-layer contradiction as a concrete, reproducible attack workflow in deployed tools, nor provides a systematic protocol for detecting such contradictions. Our work addresses this gap by instantiating provenance-laundering workflows that produce authenticated fakes, and evaluating a cross-layer audit protocol that jointly reasons over provenance metadata and pixel-level watermark signals.
@@ 3 Threat Model
The adversary we consider can generate AI images using publicly available models, embed or preserve watermarks using open-source tools, and pass images through C2PA-compliant signing workflows to attach valid manifests. The adversary cannot forge cryptographic signatures, compromise certificate authorities, or break standard cryptographic primitives. The attack is therefore not cryptanalytic but procedural as the adversary simply omits the AI-origin assertion from the manifest, which is permitted by the specification since C2PA does not mandate that signers declare generative origins [11]. While the Integrity Clash can arise through other paths such as soft-binding collisions, this work focuses on provenance laundering, in which watermarked AI-generated content is passed through a legitimate signing workflow without disclosure of its synthetic origin, as the most operationally straightforward vector.
@@ 4.3 C2PA Signing
Figure 2: Comparison of an honestly declared AI-generated image manifest (Q4a, top) and an authenticated fake (Q4b, bottom). The digitalSourceType field is entirely absent in the attack manifest, and the action and software agent are replaced with generic editing descriptors. The attack requires no cryptographic compromise, only the semantic omission of the AI origin assertion, which the C2PA specification does not mandate [11].
@@ 4.5 Cross-Layer Audit Protocol
Current verification workflows evaluate provenance metadata and watermark signals independently. A C2PA verifier validates the manifest signature and displays the declared assertions; a watermark detector decodes the pixel-level payload and reports detection status. Neither conditions on the output of the other, which means the contradictory state defined as Q4b in Section 3.1 is invisible to both. The audit protocol proposed here closes this gap by jointly evaluating both layers and applying consistency rules derived from the conflict matrix.
@@ 5.1 Core Pipeline
Figure 3 illustrates this result against deployed verification infrastructure. We submitted the same AI-generated, watermarked image to the Content Credentials Verify [^2] https://verify.contentauthenticity.org tool under three conditions: with no manifest attached, with the honest AI-disclosure manifest, and with the misleading human-edited manifest. The verifier’s output changes entirely based on the attached metadata. When no manifest is present, the tool reports no content credential. When the honest manifest is attached, the tool correctly identifies the image as AI-generated and surfaces the generative model. When the misleading manifest is attached, the tool displays the image as human-edited with no mention of AI involvement. In both signed cases, the only warning raised concerns the certificate issuer, which is unrecognized because our research certificate is not in the platform’s trust store. The verifier has no mechanism to inspect or cross-reference the pixel-level watermark signal. A CA-issued certificate from a C2PA-compliant tool would eliminate the issuer warning while the underlying semantic contradiction remains undetected. We note that Q3 (Authenticated Content), in which a valid manifest is present but no watermark is detected, is not instantiated as a separate pipeline. The baseline images confirm the detector’s behavior on unwatermarked content, and Q3 would require either a human-captured image signed with C2PA or the removal of the watermark before signing, neither of which is the attack vector under investigation.
@@ 5.2 Cross-Layer Audit Protocol
Table 3 reports classification accuracy across all conflict-matrix states. The protocol correctly classifies every image in every condition, including all 2,000 Q4b instances spanning the unperturbed and three perturbed variants. No image from the non-attack pipelines is misclassified as an authenticated fake, and no authenticated fake evades detection. The protocol requires no infrastructure beyond what already exists in isolation, a C2PA manifest verifier and a watermark detector. Its contribution is the joint evaluation logic that cross-references the two outputs and flags semantic inconsistencies that neither layer surfaces independently. We note that the perfect classification reported here reflects the strong robustness of Pixel Seal under the perturbations tested. A weaker watermarking scheme or more aggressive degradation could push images below the detection threshold, at which point the misleading manifest becomes the sole remaining signal with no contradicting evidence, a failure mode that is arguably more dangerous than the Integrity Clash itself.
Figure 4: Bit accuracy distributions across experimental conditions ($N=500$ per condition). The dashed line marks the 0.75 detection threshold. Baseline images (no watermark) cluster at chance level ($\sim$0.50), confirming no false detections. All watermarked conditions remain well above the threshold, with screenshot simulation producing the widest spread (min==0.906). No image crosses the detection boundary under any perturbation.
@@ 6.2 Limitations & Future Work
The Integrity Clash surfaces a structural condition in which two independent verification infrastructures produce irreconcilable judgments about the same asset. This has implications beyond detection failure. When a valid C2PA manifest attributes an image to a human editor while an embedded watermark identifies it as the output of a generative model, the result is not merely an inconsistency but a dispute between competing provenance claims, each backed by a different trust infrastructure, with no established mechanism to determine which takes precedence. As adoption of both systems accelerates, the frequency of such disputes will increase, and the absence of cross-layer adjudication will become a practical liability for any platform, newsroom, or legal proceeding that relies on either signal in isolation.

### v2

# Authenticated Contradictions from Desynchronized Provenance and Watermarking

@@ Abstract
Cryptographic provenance standards such as C2PA and invisible watermarking are positioned as complementary defenses for content authentication, yet the two verification layers are technically independent: neither conditions on the output of the other. This work formalizes and empirically demonstrates the Integrity Clash, a condition in which a digital asset carries a cryptographically valid C2PA manifest asserting human authorship while its pixels simultaneously carry a watermark identifying it as AI-generated, with both signals passing their respective verification checks in isolation. We construct metadata washing workflows that produce these authenticated fakes through standard editing pipelines, requiring no cryptographic compromise, only the semantic omission of a single assertion field permitted by the current C2PA specification. To close this gap, we propose a cross-layer audit protocol that jointly evaluates provenance metadata and watermark detection status, achieving 100% classification accuracy across 3,500 test images spanning four conflict-matrix states and three realistic perturbation conditions. Our results demonstrate that the gap between these verification layers is unnecessary and technically straightforward to close. Our code is available at https://github.com/ANCP2021/integrity-clash.
@@ 1 Introduction
This framing, however, obscures a structural vulnerability that follows directly from the architecture of both systems. The two verification layers are technically independent where C2PA manifest validation is a cryptographic signature check over metadata [8, 7, 45], while watermark detection is a signal-decoding operation over pixel data [54, 55]. Neither procedure conditions on the output of the other. As a result, a digital asset can carry a cryptographically valid C2PA manifest asserting one provenance history while its pixels simultaneously carry a watermark signal that contradicts that history, with both signals passing their respective verification checks in isolation. In this paper, we formalize and empirically demonstrate this condition, which we refer to as the Integrity Clash. Concretely, an AI-generated image can be passed through a C2PA-compliant editing tool, signed with valid credentials that assert only a human editing action, and emerge with a valid provenance manifest while still containing a watermark that identifies it as synthetically generated. A verifier that inspects only one layer receives an incomplete account of the asset’s origin.
@@
Prior work has examined each verification layer in isolation. Research on watermarking evaluates embedding robustness under compression, cropping, and adversarial attack [4, 53, 43], while research on provenance standards analyzes manifest integrity, soft-binding collision risks, and metadata stripping vulnerabilities [7, 9]. Recent empirical audits have documented fragmented adoption of these systems across generative AI platforms [37, 42], and preliminary frameworks have proposed multi-signal consistency checks as a forensic methodology [45, 23]. However, no existing work empirically demonstrates the cross-layer contradiction as a concrete, reproducible vulnerability in deployed pipelines, nor provides a systematic protocol for detecting such contradictions in deployed pipelines. This paper addresses that gap. We make three contributions:
@@ 2.2 Invisible Watermarking for Generative AI
Recent work has shown that invisible watermarks are not immutable. Benchmarks such as WAVES [4] evaluate robustness under compression, rescaling, and noise, and adversarial studies have demonstrated practical removal and transplantation attacks against open-source watermarking tools [31, 35]. The winning system in the NeurIPS 2024 Invisible Watermark Removal Challenge [44, 15], for example, uses generative-model-based regeneration to strip a variety of watermark signals while maintaining high perceptual quality, and Zhao et al. [53] prove that a broad class of invisible image watermarks are removable using generative AI. These findings establish that watermark signals can be weakened under determined attackers, but also confirm that under standard editorial workflows, watermarks remain persistent to produce detectable cross-layer inconsistencies with co-existing metadata.
@@ 2.3 Related Work
Ecosystem audits have begun to document the gap between these layers. Rijsbosch et al. [42] report that only 38% of surveyed AI image generators implement adequate watermarking and just 18% integrate with C2PA-style labeling standards, establishing that the fragmentation enabling cross-layer inconsistencies is empirically widespread. Nemecek et al. [37] argue that without independent verification across technical, audit, and enforcement layers, watermarking serves as symbolic compliance rather than effective oversight. Simmons and Winograd [45] define success and failure scenarios when provenance and watermark signals co-exist in broadcast media, while agent-based forensic pipelines have been explored for multimodal detection [27]. These contributions identify the fragmentation problem or propose frameworks, but none demonstrate the cross-layer contradiction as a reproducible specification-level vulnerability or evaluate a detection protocol against it.
Industry and standards-focused analyses reinforce this picture at a systems level. Recent work on media integrity [51] concludes that high-confidence validation requires aligning multiple signals rather than trusting any single mechanism in isolation, and the Origin Lens framework [30] advocates combined provenance verification and AI detection but does not examine cases where a valid manifest coexists with a contradictory watermark. None, to our knowledge, empirically demonstrates the cross-layer contradiction as a concrete, reproducible vulnerability in deployed tools, nor provides a systematic protocol for detecting such contradictions. Our work addresses this gap by constructing workflows that exploit this specification-level omission to produce authenticated fakes, and evaluating a cross-layer audit protocol that jointly reasons over provenance metadata and pixel-level watermark signals.
@@ 3 Threat Model
The adversary we consider can generate AI images using publicly available models, embed or preserve watermarks using open-source tools, and pass images through C2PA-compliant signing workflows to attach valid manifests. The adversary cannot forge cryptographic signatures, compromise certificate authorities, or break standard cryptographic primitives. The resulting vulnerability is therefore not cryptanalytic but procedural, as the adversary simply omits the AI-origin assertion from the manifest, which is permitted by the specification since C2PA does not mandate that signers declare generative origins [11]. While the Integrity Clash can arise through other paths such as soft-binding collisions, this work focuses on provenance laundering, in which watermarked AI-generated content is passed through a legitimate signing workflow without disclosure of its synthetic origin, as the most operationally straightforward vector.
@@ 4.3 C2PA Signing
Figure 2: Comparison of an honestly declared AI-generated image manifest (Q4a, top) and an authenticated fake (Q4b, bottom). The digitalSourceType field is entirely absent in the misleading manifest, and the action and software agent are replaced with generic editing descriptors. The workflow requires no cryptographic compromise, only the semantic omission of the AI origin assertion, which the C2PA specification does not mandate [11].
@@ 4.5 Cross-Layer Audit Protocol
Because existing verification workflows evaluate provenance metadata and watermark signals independently (Section 3), the contradictory state Q4b is invisible to both. The audit protocol proposed here closes this gap by jointly evaluating both layers and applying consistency rules derived from the conflict matrix.
@@ 5.1 Core Pipeline
Figure 3 illustrates this result against deployed verification infrastructure. We submitted the same AI-generated, watermarked image to the Content Credentials Verify [^2] https://verify.contentauthenticity.org tool under three conditions: with no manifest attached, with the honest AI-disclosure manifest, and with the misleading human-edited manifest. The verifier’s output changes entirely based on the attached metadata. When no manifest is present, the tool reports no content credential. When the honest manifest is attached, the tool correctly identifies the image as AI-generated and surfaces the generative model. When the misleading manifest is attached, the tool displays the image as human-edited with no mention of AI involvement. In both signed cases, the only warning raised concerns the certificate issuer, which is unrecognized because our research certificate is not in the platform’s trust store. The verifier has no mechanism to inspect or cross-reference the pixel-level watermark signal. A CA-issued certificate from a C2PA-compliant tool would eliminate the issuer warning while the underlying semantic contradiction remains undetected. We note that Q3 (Authenticated Content), in which a valid manifest is present but no watermark is detected, is not instantiated as a separate pipeline. The baseline images confirm the detector’s behavior on unwatermarked content, and Q3 would require either a human-captured image signed with C2PA or the removal of the watermark before signing, neither of which is the vulnerability under investigation.
@@ 5.2 Cross-Layer Audit Protocol
Table 3 reports classification accuracy across all conflict-matrix states. The protocol correctly classifies every image in every condition, including all 2,000 Q4b instances spanning the unperturbed and three perturbed variants. No image from the non-adversarial pipelines is misclassified as an authenticated fake, and no authenticated fake evades detection. The protocol requires no infrastructure beyond what already exists in isolation, a C2PA manifest verifier and a watermark detector. Its contribution is the joint evaluation logic that cross-references the two outputs and flags semantic inconsistencies that neither layer surfaces independently. We note that the perfect classification reported here reflects the strong robustness of Pixel Seal under the perturbations tested; the implications of weaker watermarking schemes are discussed in Section 6.2.
Figure 4: Bit accuracy distributions across experimental conditions (NN==500 per condition). The dashed line marks the 0.75 detection threshold. Baseline images (no watermark) cluster at chance level ($\sim$0.50), confirming no false detections. All watermarked conditions remain well above the threshold, with screenshot simulation producing the widest spread (min==0.906). No image crosses the detection boundary under any perturbation.
@@ 6.1 Implications for Deployed Verification Infrastructure
The cross-layer audit protocol evaluates only C2PA assertions and watermark status, but its consistency-checking logic is extensible to additional signals. Passive forensic methods such as frequency-domain analysis and learned AI-generation classifiers provide independent evidence of synthetic origin that depends on neither a declared manifest nor an embedded watermark. Integrating these as supplementary inputs would strengthen detection when one primary layer is absent or degraded, and agent-based forensic pipelines [27] and multi-signal consistency frameworks [23, 30] offer architectural models for doing so.
@@ 6.2 Limitations & Future Work
As adoption of both systems accelerates, the absence of cross-layer adjudication will become a practical liability for any platform, newsroom, or legal proceeding that relies on either signal in isolation.
@@ 7 Conclusion
Acknowledgments
This work was supported in part by the NSF under CNS-2247795, 2427505, and 2141622, by the Office of Naval Research under N00014-22-1-2680, and by a gift from Cisco.